### v1

@@ 3.5 Loss Function
$L_{M}=\frac{1}{N}\sum_{i=1}^{N}\left(\widehat{y_{i}}-y_{i}\right)^{2}$ (19)
@@
$l_{j}(y,\hat{y})=\frac{2\mu_{y}\mu_{\hat{y}}+C_{1}}{\mu_{y}^{2}+\mu_{\hat{y}}^{2}+C_{1}}$ (20)
$c_{j}(y,\hat{y})=\frac{2\sigma_{y}\sigma_{\hat{y}}+C_{2}}{\sigma_{y}^{2}+\sigma_{\hat{y}}^{2}+C_{2}}$ (21)
$s_{j}(y,\hat{y})=\frac{\sigma_{y\hat{y}}+C_{3}}{\sigma_{y}\sigma_{\hat{y}}+C_{3}}$ (22)
@@
$\operatorname{MS-SSIM}=\left[l_{j}^{\alpha}\cdot\prod_{j=1}^{L}c_{j}^{\beta}s_{j}^{\gamma}\right]$ (23)
@@
$L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}$ (25)
@@ 4 Experimental Results and Discussions
4.1 Dataset
A significant barrier to progress in ground-based cloud extrapolation has been the lack of high-quality, large-scale public datasets. The performance of deep learning models is fundamentally tied to the data they are trained on, yet existing benchmarks suffer from critical limitations that fail to capture the true complexity of cloud dynamics. For instance, the popular TSISD dataset [STANet] features a low spatial resolution of only (224 $\times$ 224 $\times$ 3 for each image) and contains significant visual occlusions from camera hardware, which can introduce artifacts and degrade prediction precision.
To address this gap and provide a more challenging and realistic benchmark for the research community, we developed and publicly release the ASI-Cloud Image Sequence (ASI-CIS) dataset, a core contribution of this paper. In this dataset, 1, 400 sequences are used for training and 700 sequences are used for testing. To eliminate data similarity between splits, the training and testing sets are acquired from temporally distinct periods, with validation performed via five-fold cross-validation. The size of each image is 512 $\times$ 512 pixels, and consecutive frames are captured at 30-second intervals using the All Sky Smage (ASI-DC-TK02). This device is located in Xiqing District, Tianjin, China (geographic coordinates: 117.03°E, 39.10°N). Data collection spans multiple seasons, with acquisition times ranging from 08:00 to 17:00 local time, ensuring robust coverage of diverse meteorological conditions for ground-based cloud image sequence extrapolation tasks. Table 1 summarizes the ASI-CIS dataset according to various weather conditions and quantities. The dataset includes samples under diverse weather conditions, including sunny, cloudy/rainy scenarios, with 1421 and 579 sequences, respectively.
@@ 4.1 Dataset
The observed data imbalance primarily stems from the temperate monsoon climate in the acquisition region, where clear-sky conditions occur significantly more frequently than cloudy/rainy weather. This inherent meteorological distribution directly induces substantial disparities in raw data collection volumes. Furthermore, suboptimal acquisition conditions during precipitation events, which are characterized by heavy rainfall, low illumination, and potential lens contamination, frequently yield poor-quality imagery. Cloudy conditions present additional challenges due to complex cloud structures (e.g., stratocumulus-cumulus mixtures) that exceed preset sensor parameters, rendering cloud formations unrecognizable. Representative samples across weather conditions are shown in Fig. 7.
@@ 4.2 Evaluation Metrics
Round-based remote sensing cloud image sequence extrapolation task is essentially an image tracking task. Therefore, to evaluate the performance of our proposed method, several common metrics are employed, including the Mean Squared Error (MSE), Structural Similarity Index (SSIM), and Peak Signal-to-Noise Ratio (PSNR).
@@
$\mathrm{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}(I(i,j)-K(i,j))^{2}$ (26)
where $I$ and $K$ denote the predicted and GT images, respectively, and $m\times n$ denotes the image dimensions.
@@
$\operatorname{SSIM}(I,K)=\frac{\left(2\mu_{I}\mu_{K}+c_{1}\right)\left(2\sigma_{IK}+c_{2}\right)}{\left(\mu_{I}^{2}+\mu_{K}^{2}+c_{1}\right)\left(\sigma_{I}^{2}+\sigma_{K}^{2}+c_{2}\right)}$ (27)
@@
$\operatorname{PSNR}=10\cdot\log_{10}\left(\frac{\mathrm{MAX}_{I}^{2}}{\mathrm{MSE}}\right)$ (28)
@@ 4.3 Implementation Details
The code of our proposed USF-Net is built on the PyTorch framework with the Python programming language. The experimental platform and environment are as follows: a computer with Ubuntu 18.04, an Intel (R) Xeon (R) Gold 5318Y CPU @ 2.10 GHz, and two NVIDIA A40 GPUs with graphics memory of 48 GB. During the training phase, we set the initial learning rate to $10^{-3}$. To ensure the smooth training of the network, we used the SGD optimizer with 0.9 momentum and $10^{-4}$ weight decay. The training batch size is set to 4, and the learning rate will decay every 10 epochs until it reaches the minimum learning rate of $10^{-5}$. Consistent with the previous study [STANet], we employ the scheduled sampling strategy to mitigate generalization errors and enhance the ability to learn long-term spatiotemporal dynamics. Specifically, the occlusion rate $P$ applied to input images is progressively increased from 0 to 1 during the training duration. The proposed USF-Net is trained for 1,00 epochs on our dataset, and the training time was 4.8 h.
@@ 5 Conclusion
Accurate and efficient extrapolation of ground-based cloud image sequences is a critical enabling technology for the stable integration of photovoltaic power systems. In this paper, we addressed key limitations in existing deep learning methods by proposing USF-Net, a novel framework for cloud image sequence extrapolation that introduces spatiotemporal architecture to unify the modeling of spatial and temporal information a novel framework that unifies the modeling of spatial and temporal information. Our contributions are threefold. First, we introduced a unified spatiotemporal architecture where temporal flow information explicitly guides spatial feature learning via a novel Temporal Guidance Module (TGM). Second, we designed a Unified Spatiotemoral Module (USTM) that contains a Spatial Selection Module (SSM) to dynamically capture multi-scale cloud context and a Temporal Agent Attention Module (TAM) to model long-range temporal dependencies with linear complexity efficiently. Third, we developed a Dynamic Update Module (DUM) in the decoder that leverages initial temporal states to effectively mitigate the “ghosting effect” and preserve motion fidelity. Extensive experiments on our newly proposed high-resolution ASI-CIS dataset demonstrate that USF-Net significantly outperforms existing SOTA methods. Furthermore, ablation studies rigorously validate the effectiveness of each of our proposed modules. The results confirm that USF-Net establishes a new benchmark for the task, achieving superior prediction accuracy while maintaining high computational efficiency. Future work will focus on extending USF-Net to additional photovoltaic power prediction tasks, further enhancing the method’s real-time inference capability.
CRediT authorship contribution statement
Penghui Niu: Writing - original draft, Writing – review &\& editing, Visualization, Validation, Software, Methodology, Investigation, Funding acquisition. Taotao Cai: Writing – review &\& editing, Methodology, Investigation, Formal analysis. Jiashuai She: Writing - original draft, Writing – review &\& editing, Validation, Data curation. Yajuan Zhang: Resources, Project administration, Investigation. Junhua Gu: Supervision, Resources, Project administration, Funding acquisition, Data curation. Ping Zhang: Resources, Project administration, Funding acquisition. Jungong Han: Writing – review &\& editing, Supervision, Methodology. Jianxin Li: Writing – review &\& editing, Supervision, Project administration, Methodology.
Declaration of competing interest
The authors declare that they have no known competing finan cial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The proposed ASI-CIS dataset and source code will be available at https://github.com/she1110/ASI-CIS.
Acknowledgments
The research was supported by the National Nature Science Foundation of China under Grant No. 62206085, supported by the Innovation Capacity Improvement Plan Project of Hebei Province under 22567603H, and supported by the Interdisciplinary Postgraduate Training Program of Hebei University of Technology under HEBUT-Y-XKJC-2022101.

### v2

@@ 3.5 Loss Function
$L_{M}=\frac{1}{N}\sum_{i=1}^{N}\left(\widehat{y_{i}}-y_{i}\right)^{2}.$ (19)
@@
[…]
@@
$\operatorname{MS-SSIM}=\left[l_{j}^{\alpha}\cdot\prod_{j=1}^{L}c_{j}^{\beta}s_{j}^{\gamma}\right].$ (23)
@@
$L=\lambda_{1}L_{M}+\lambda_{2}L_{MS}+\lambda_{3}L_{C}.$ (25)
@@ 4 Experimental Results and Discussions
4.1 Dataset
The scarcity of high-quality, large-scale public datasets remains a primary bottleneck in ground-based cloud extrapolation research. The predictive capability of deep learning models is intrinsically linked to the spatiotemporal fidelity of their training data. Existing benchmarks, such as the TSISD dataset [STANet], utilize a constrained spatial resolution of $224\times 224$ pixels and frequently exhibit visual occlusions arising from sensor hardware. These limitations introduce aliasing artifacts and impede the precise modeling of complex cloud dynamics.
@@ 4.1 Dataset
To address these challenges and provide a robust benchmark for the remote sensing community, we introduce the ASI-Cloud Image Sequence (ASI-CIS) dataset. The ASI-CIS dataset was acquired using an All-Sky Imager (ASI-DC-TK02) stationed at the meteorological observation facility in Xiqing District, Tianjin, China (geographic coordinates: $117.03^{\circ}$E, $39.10^{\circ}$N). The acquisition device features a fish-eye lens providing a hemispherical field of view, housed within a weatherproof enclosure to ensure operational continuity. The dataset consists of high-resolution images ($512\times 512$ pixels) captured at fixed 30-second intervals. This temporal granularity is critical for capturing rapid cloud deformations and velocity changes characteristic of the lower troposphere. The data collection campaign spanned multiple seasons with daily acquisition windows from 08:00 to 17:00 local time, ensuring coverage across a diverse spectrum of solar angles and illumination conditions.
@@
The dataset comprises a total of 2,100 sequences, partitioned into 1,400 sequences for training and 700 sequences for testing. To prevent temporal data leakage, where the model inadvertently learns from future frames closely correlated with the test set, the training and testing partitions were acquired from temporally distinct periods. Validation was conducted using a rigorous five-fold cross-validation protocol. Table 1 summarizes the ASI-CIS dataset according to various weather conditions and quantities. Specifically, the dataset contains 1,421 sunny sequences compared to 579 cloudy/rainy sequences. While this imbalance presents a challenge for class-agnostic learning, it faithfully reflects the real-world operational environment of photovoltaic power plants in this climatic zone. Furthermore, data acquisition during precipitation events presents unique challenges. Heavy rainfall and low illumination can degrade image signal-to-noise ratios, while complex multi-layer cloud configurations (e.g., stratocumulus-cumulus mixtures) often exceed the dynamic range of standard sensors. The ASI-CIS dataset preserves these challenging samples to rigorously test model robustness under adverse meteorological conditions. Representative samples across weather conditions are shown in Fig. 7.
@@ 4.2 Evaluation Metrics
Ground-based remote sensing cloud image sequence extrapolation is fundamentally a spatiotemporal predictive learning task. To comprehensively evaluate the performance of USF-Net, we employ three standard metrics: Mean Squared Error (MSE), Structural Similarity Index (SSIM), and Peak Signal-to-Noise Ratio (PSNR).
@@
$\mathrm{MSE}=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}(I(i,j)-K(i,j))^{2}.$ (26)
where $I$ and $K$ denote the predicted and GT images, respectively, and $mn$ denotes the spatial dimensions of the image.
@@
$\operatorname{SSIM}(I,K)=\frac{\left(2\mu_{I}\mu_{K}+c_{1}\right)\left(2\sigma_{IK}+c_{2}\right)}{\left(\mu_{I}^{2}+\mu_{K}^{2}+c_{1}\right)\left(\sigma_{I}^{2}+\sigma_{K}^{2}+c_{2}\right)}.$ (27)
@@
$\operatorname{PSNR}=10\cdot\log_{10}\left(\frac{\mathrm{MAX}_{I}^{2}}{\mathrm{MSE}}\right).$ (28)
@@ 4.3 Implementation Details
The USF-Net framework was implemented using the PyTorch library and executed on a high-performance computing platform equipped with an Intel Xeon Gold 5318Y CPU (@ 2.10 GHz) and two NVIDIA A40 GPUs (48 GB VRAM).The network was optimized using the Stochastic Gradient Descent (SGD) algorithm with a momentum of 0.9 and a weight decay of $10^{-4}$ to prevent overfitting. The initial learning rate was set to $10^{-3}$. We employed a step-decay learning rate scheduler, where the learning rate is reduced by a factor of 10 every 10 epochs until reaching a minimum floor of $10^{-5}$. The training batch size was fixed at 4 samples per iteration.To mitigate exposure bias and enhance the model’s capability to capture long-term spatiotemporal dynamics, we adopted a scheduled sampling strategy. The sampling probability $P$, which governs the substitution of ground truth frames with model-generated predictions during the training sequence, was linearly increased from 0 to 1 over the course of the training duration. The model was trained for 100 epochs on the ASI-CIS dataset, with a total convergence time of approximately 4.8 hours.
@@ 5 Conclusion
Precise and efficient extrapolation of ground-based cloud image sequences constitutes a pivotal enabling technology for mitigating the intermittency inherent in photovoltaic power integration. This study circumvents the limitations of prevailing deep learning methodologies by introducing USF-Net, a novel framework that orchestrates a spatiotemporal architecture to synthesize spatial features with temporal dynamics. The primary technical contributions are realized through three specialized components. First, a TGM was developed to explicitly modulate spatial feature learning using temporal flow, thereby ensuring coherent spatiotemporal representations. Second, the USTM was engineered, comprising a SSM for dynamic multi-scale context capture and a TAM that resolves long-range dependencies with linear computational complexity. Third, a DUM integrated into the decoder leverages initial temporal states to mitigate the ghosting effect, preserving high-fidelity motion signatures. Extensive empirical validation on the newly curated high-resolution ASI-CIS dataset demonstrates that USF-Net surpasses current SOTA benchmarks. Concurrently, comprehensive ablation studies corroborate the individual efficacy of the proposed TGM, USTM, and DUM components. By achieving a superior equilibrium between prediction accuracy and computational efficiency, USF-Net establishes a new benchmark for the domain. Future research will explore the extensibility of USF-Net to broader photovoltaic power forecasting tasks and the further optimization of real-time inference capabilities.
[…]
Acknowledgements The research was supported by the National Natural Science Foundation of China under Grant No. 62206085, supported by the Innovation Capacity Improvement Plan Project of Hebei Province under 22567603H, and supported by the Interdisciplinary Postgraduate Training Program of Hebei University of Technology under HEBUT-Y-XKJC-2022101.